# Hyperband-based Bayesian Optimization for Black-box Prompt Selection

### v1

@@ E.1 Main Results
  • TRIPLE-SH vs. HbBoPs, $z=1.49,p=1.00$
  • TRIPLE=GSE vs. HbBoPs, $z=1.34,p=1.00$
Conclusions are mostly similar to the analysis with respect to validation performance, however, while HbBoPs improves over TRIPLE-SH and TRIPLE-GSE also with respect to test performance, the effects are not strong enough to be considered statistically significant at the $\alpha=0.01$ level. For brevity, we do not include results for fractions of $0.25$ or $0.50$ here, where we observed similar results but HbBoPs again more strongly outperforming the other methods.
Table 9: Normalized test error of each method on each benchmark for Claude 3 Haiku. Averaged over repetitions. Standard errors in parentheses. For visual analysis, we highlight all methods that have a mean performance that is less or equal to the mean performance of the best method plus two times its standard error.
@@
Table 10: Normalized test error of each method on each benchmark for LLAMA3 8B Instruct. Averaged over repetitions. Standard errors in parentheses. For visual analysis, we highlight all methods that have a mean performance that is less or equal to the mean performance of the best method plus two times its standard error.
@@
Table 11: Normalized test error of each method on each benchmark for Mistral 7B Instruct. Averaged over repetitions. Standard errors in parentheses. For visual analysis, we highlight all methods that have a mean performance that is less or equal to the mean performance of the best method plus two times its standard error.
@@ E.2 Ablation Study
Here we provide additional analyses of the ablation results reported in Section 5.3. To test the significance of each component on the validation and test performance of HbBoPs, we again conduct a linear mixed effects model analysis. We model the unaggregated performance as a function of the fraction of LLM calls (i.e., over time; starting after the initial design of full-fidelity methods, i.e., after a fraction of $0.40$) and random intercepts for each benchmark scenario (benchmark task and LLM combination). To test the global hypothesis that there is an effect of the components on performance, we test an intercept model against a model including an effect of the factor method. If we reject the null hypothesis we proceed with a Tukey post hoc test (corrected for multiple testing) to perform pairwise comparisons. We test at the $\alpha=0.01$ level.
Looking at the validation error, we reject the global null hypothesis of no effect of the method on the anytime performance ($\chi^{2}(5)=35184.28,p<1\mathrm{e}{-4}$). The relevant pairwise comparison results are the following:
@@
  • BoPs (structural-aware DK-GP vs. HB, $z=52.92,p<1\mathrm{e}{-4}$
@@
We can conclude that using a DK-GP significantly improves over vanilla BO, that a structural-aware DK-GP improves over the non structural-aware DK-GP, that HB further improves of the structural-aware DK-GP and HbBoPs further improving over HB.
We perform the same analysis for the test performance and reject the global hypothesis of no effect ($\chi^{2}(5)=14578.73,p<1\mathrm{e}{-4}$). The relevant pairwise comparison results are the following:
@@
Conclusions are the same as when looking at validation performance.
@@ E.3 Encoder Sensitivity
Here we provide additional analyses of the encoder sensitivity results reported in Section 5.4.
@@
To test whether the choice of encoder model does make a difference for the final performance of HbBoPs, we again conduct a linear mixed effects model analysis. We model the unaggregated performance at a fraction of $1.00$ total LLM calls involving random intercepts for each benchmark scenario (benchmark task and LLM combination). To test the global hypothesis that there is an effect of the encoder on performance, we test an intercept model against a model including an effect of the factor encoder. We test at the $\alpha=0.01$ level. For the both the validation and test performance, we can not reject the null hypothesis of the encoder making no difference, $\chi^{2}(2)=4.69,p=0.096$ and $\chi^{2}(2)=3.85,p=0.146$ respectively.
E.4 An Ablation of our Design Decisions for HB
As mentioned in Section 3 and Appendix C, adapting HB to prompt selection involves several design decisions. Here we provide an ablation of these decisions.
@@ E.4 An Ablation of our Design Decisions for HB
While vanilla HB for Hyperparameter optimization would return the configuration with the lowest validation error as the (anytime) incumbent, this is not sensible for prompt selection as the fidelity directly influences the noise of the validation error. We therefore always return the prompt with the lowest validation error among all prompts that have been evaluated on the (current) highest fidelity level. To ablate this design decision, we run HB for prompt selection with this incumbent selection mechanism and compare to the incumbent selection mechanism that simply selects the prompt with the lowest validation error. The experimental setup is exactly the same as for the results reported in the main paper. We visualize the (oracle) normalized validation and test error of the best prompt found by HB under each incumbent selection scheme in Figure 6. As before, for visualization purposes the validation error of the incumbent is computed here in an oracle setting (i.e., using all validation instances), whereas during the selection process the anytime incumbent itself was selected based on its validation error computed on fewer validation instances. Looking at the validation error (Figure 6(a)) of HB as used by us (selecting the incumbent as the prompt with the lowest validation error among all prompts evaluated on the highest fidelity level), we can see that validation error keeps decreasing as optimization processes. In contrast, if we would perform the incumbent selection simply by choosing the prompt with the lowest validation error (ignoring the fidelity level), as in HB (incumbent lowest validation error), we can see that optimization progress quickly stagnates. This is due to the incumbent no longer being updated as on the initially lower fidelity levels, lower validation errors can be observed due to noisy performance estimates based on few validation instances. Similar conclusions hold for the test error (Figure 6(b)).
@@
Figure 6: Anytime normalized error of the best prompt found by each HB variant, averaged over different benchmark scenarios. Lower is better. Log scale. Ribbons represent standard errors.
@@
To test whether the choice of selection mechanism model does make a difference for the final performance of HB, we again conduct a linear mixed effects model analysis. We model the unaggregated performance at a fraction of $1.00$ total LLM calls involving random intercepts for each benchmark scenario (benchmark task and LLM combination). To test the global hypothesis that there is an effect of the incumbent selection mechanism on performance, we test an intercept model against a model including an effect of the factor selection mechanism. We test at the $\alpha=0.01$ level. For both validation and test performance we reject the null hypothesis of no effect of the selection mechanism $\chi^{2}(1)=448.05,p<1\mathrm{e}{-4}$ and $\chi^{2}(1)=164.17,p<1\mathrm{e}{-4}$ respectively.
@@
Another design decision for adapting HB to prompt selection is concerned whether prompts should be evaluated on the same, random validation instances within a stage of a given budget of if they should be evaluated on their own random samples. The final related design decision involves whether validation instances of higher stages for a given bracket are constructed to be supersets of the validation instances used in lower stages (as described in Appendix C) which allows for further speed-ups due to caching.
@@
To investigate the effect of using the same random vs. truly random instances for each prompt and the effect of validation instances used in higher stages of a bracket being supersets of the validation instances used in lower stages, we run HB for prompt selection varying these two components. The experimental setup is exactly the same as for the results reported in the main paper. We visualize the (oracle) normalized validation and test error of the best prompt found by HB under each incumbent selection scheme in Figure 7. As before, for visualization purposes the validation error of the incumbent is computed here in an oracle setting (i.e., using all validation instances), whereas during the selection process the anytime incumbent itself was selected based on its validation error computed on fewer validation instances. Looking at the validation error (Figure 7(a)) of HB as used by us (same random instances and supersets), we can see that this variant performs best. If we would use truly random instances (but keep the superset structure) as in HB (random instances for each prompt), performance is slightly worse. Giving up the superset structure (HB (no supersets, same instances for each prompt) and HB (no supersets, random instances for each prompt) we can see that performance is substantially worse, and even more when using truly random instances for each prompt. In general we can conclude that the effect of using supersets for higher stages within a given bracket boosts the validation performance of HB. Moreover, using truly random validation instances for each prompt instead of using the same random validation instances for all prompt evaluations in a stage (i.e., the paired setting) generally worsens performance. Looking at the test performance, these statements generalize (Figure 7(b)).
@@
Figure 7: Anytime normalized error of the best prompt found by each HB variant, averaged over different benchmark scenarios. Lower is better. Log scale. Ribbons represent standard errors.
To test whether the design choices of using supersets and using the same random validation instances for each prompt evaluation within a stage make a difference for the final performance of HB, we again conduct a linear mixed effects model analysis. We model the unaggregated performance at a fraction of $1.00$ total LLM calls involving random intercepts for each benchmark scenario (benchmark task and LLM combination). To test the global hypothesis that there is an effect of the design choices on performance, we test an intercept model against a model including an effect of the factor method. We test at the $\alpha=0.01$ level.
For the validation performance at a fraction of $1.00$, we reject the global null hypothesis of no effect of methods ($\chi^{2}(3)=730.85,p<1\mathrm{e}{-4}$). The pairwise results are:
@@
Looking at test performance at a fraction of $1.00$ we also reject the global null hypothesis of no effect of methods ($\chi^{2}(3)=207.12,p<1\mathrm{e}{-4}$). The pairwise results are:
@@
  • HB (no supersets, random instances for each prompt) vs. HB, $z=11.022,p<1\mathrm{e}{-4}$
@@
Summarizing, we have seen that our design decisions made to adapt HB to prompt selection are reasonable: (1) The incumbent should be selected as the prompt with the lowest validation error among all prompts evaluated on the highest fidelity level. (2) Validation instances used to evaluate prompts in higher stages of a given bracket should be supersets of the validation instances used in lower stages. (3) Using the same (random) validation instances to evaluate prompts in each stage in general is beneficial compared to using truly random validation instances for each prompt albeit this effect is comparably small.

### v2

@@ E.1 Main Results
  • TRIPLE-SH vs. HbBoPs, $z=1.49,p=1.000$
  • TRIPLE-GSE vs. HbBoPs, $z=1.34,p=1.000$
Conclusions are largely consistent with the analysis with respect to validation error, however, while HbBoPs improves over TRIPLE-SH and TRIPLE-GSE also with respect to test error, the effects are not strong enough to be considered statistically significant at the $\alpha=0.01$ level. For brevity, we do not include results for fractions of $0.25$ or $0.50$ here, where we observed similar results but HbBoPs again more strongly outperforming the other methods.
Table 9: Normalized test error of each method on each benchmark for Claude 3 Haiku. Averaged over repetitions. Standard errors in parentheses. For visual analysis, we highlight all methods that have a mean error that is less or equal to the mean error of the best method plus two times its standard error.
@@
Table 10: Normalized test error of each method on each benchmark for LLAMA3 8B Instruct. Averaged over repetitions. Standard errors in parentheses. For visual analysis, we highlight all methods that have a mean error that is less or equal to the mean error of the best method plus two times its standard error.
@@
Table 11: Normalized test error of each method on each benchmark for Mistral 7B Instruct. Averaged over repetitions. Standard errors in parentheses. For visual analysis, we highlight all methods that have a mean error that is less or equal to the mean error of the best method plus two times its standard error.
@@ E.2 Ablation Study
Here, we provide additional analyses of the ablation results reported in Section 5.3. To test the significance of each component on the validation and test error of HbBoPs, we again conduct a linear mixed effects model analysis. We model the unaggregated performance as a function of the fraction of LLM calls (i.e., over time; starting after the initial design of full-fidelity methods, i.e., after a fraction of $0.40$) and include random intercepts for each benchmark scenario (benchmark task and LLM combination). To test the global hypothesis that there is an effect of the components on performance, we test an intercept model against a model including an effect of the factor method (corresponding to an ablation variant). If we reject the null hypothesis, we proceed with a Tukey post hoc test (corrected for multiple testing) to perform pairwise comparisons. We test at the conservative $\alpha=0.01$ level.
Examining the validation error, we reject the global null hypothesis of no effect of the method on the anytime performance ($\chi^{2}(5)=35184.28,p<1\mathrm{e}{-4}$). The relevant pairwise comparison results state as follows:
@@
  • BoPs (structural-aware DK-GP) vs. HB, $z=52.92,p<1\mathrm{e}{-4}$
@@
We can conclude that using a DK-GP significantly improves over vanilla BO, that a structural-aware DK-GP improves over the non structural-aware DK-GP, that HB improves over the structural-aware DK-GP and that HbBoPs further improves over HB.
We perform the same analysis for the test error and reject the global hypothesis of no effect ($\chi^{2}(5)=14578.73,p<1\mathrm{e}{-4}$). The relevant pairwise comparison results state as follows:
@@
Conclusions are the same as for the validation error.
@@ E.3 Encoder Sensitivity
Here, we provide additional analyses of the encoder sensitivity results reported in Section 5.4. To test whether the choice of encoder model affects the final performance of HbBoPs, we again conduct a linear mixed effects model analysis. We model the unaggregated performance at a fraction of $1.00$ total LLM calls involving random intercepts for each benchmark scenario (benchmark task and LLM combination). To test the global hypothesis that there is an effect of the encoder on performance, we test an intercept model against a model including an effect of the factor encoder. We test at the conservative $\alpha=0.01$ level. For both the validation and test error, we cannot reject the null hypothesis of the encoder making no difference, $\chi^{2}(2)=4.69,p=0.096$ and $\chi^{2}(2)=3.85,p=0.146$ respectively. We therefore conclude that HbBoPs is robust to the choice of encoder model.
E.4 Hyperband Design Choices
@@
As mentioned in Section 3 and Appendix C, adapting HB to prompt selection involves several design decisions. Here, we provide an ablation of these decisions. While vanilla HB for hyperparameter optimization would return the configuration with the lowest validation error as the (anytime) incumbent, this is not sensible for prompt selection as the fidelity directly influences the noise of the validation error. We therefore always return the prompt with the lowest validation error among all prompts that have been evaluated on the (current) highest fidelity level. To analyze the effect of this design decision, we run HB for prompt selection with this incumbent selection mechanism and compare to the incumbent selection mechanism that simply selects the prompt with the lowest validation error. The experimental setup is exactly the same as for the results reported in the main paper. We visualize the (oracle) normalized validation and test error of the best prompt found by HB under each incumbent selection scheme in Figure 6. As before, for visualization purposes the validation error of the incumbent is computed here in an oracle setting (i.e., using all validation instances), whereas during the selection process the anytime incumbent itself was selected based on its validation error computed on fewer validation instances. Examining the validation error (Figure 6(a)) of HB as used by us (selecting the incumbent as the prompt with the lowest validation error among all prompts evaluated on the highest fidelity level), we observe that the validation error of the incumbent keeps decreasing as optimization progresses. In contrast, if we would perform the incumbent selection simply by choosing the prompt with the lowest validation error (ignoring the fidelity level), as in HB (incumbent lowest validation error), we observe that optimization progress stagnates quickly. This occurs because the incumbent is no longer updated, since, at lower fidelity levels, noisy performance estimates can result in artificially low validation errors. Similar conclusions hold for the test error (Figure 6(b)).
@@
Figure 6: Normalized error (log scale) of the best prompt found by each HB incumbent selection mechanism, averaged over benchmarks. Lower is better. Ribbons represent SE.
@@
To test whether the choice of selection mechanism does make a difference for the final performance of HB, we again conduct a linear mixed effects model analysis. We model the unaggregated performance at a fraction of $1.00$ total LLM calls involving random intercepts for each benchmark scenario (benchmark task and LLM combination). To test the global hypothesis that there is an effect of the incumbent selection mechanism on performance, we test an intercept model against a model including an effect of the factor selection mechanism. We test at the conservative $\alpha=0.01$ level. For both validation and test error we reject the null hypothesis of no effect of the selection mechanism $\chi^{2}(1)=448.05,p<1\mathrm{e}{-4}$ and $\chi^{2}(1)=164.17,p<1\mathrm{e}{-4}$ respectively. We therefore conclude that our incumbent selection mechanism is superior.
@@
Another design decision for adapting HB to prompt selection is concerned with whether prompts should be evaluated on the same random validation instances within a stage or on their own random samples. The final related design decision involves whether validation instances of higher stages for a given bracket are constructed to be supersets of the validation instances used in lower stages (as described in Appendix C) which allows for further speed-ups due to caching.
@@
To investigate the effect of using the same random vs. truly random instances for each prompt and the effect of validation instances used in higher stages of a bracket being supersets of the validation instances used in lower stages, we run HB for prompt selection varying these two components. The experimental setup is exactly the same as for the results reported in the main paper. We visualize the (oracle) normalized validation and test error of the best prompt found by HB under each incumbent selection scheme in Figure 7. As before, for visualization purposes the validation error of the incumbent is computed here in an oracle setting (i.e., using all validation instances), whereas during the selection process the anytime incumbent itself was selected based on its validation error computed on fewer validation instances. Examining the validation error (Figure 7(a)) of our proposed HB (same random instances and supersets), we can see that this variant performs best. If we use truly random instances (but keep the superset structure) as in HB (random instances for each prompt), performance is slightly worse. Giving up the superset structure (HB (no supersets, same instances for each prompt) and HB (no supersets, random instances for each prompt) we can see that performance is substantially worse, even more so when using truly random instances for each prompt. In general, we can conclude that the effect of using supersets for higher stages within a given bracket boosts the performance of HB. Moreover, using truly random validation instances for each prompt instead of using the same random validation instances for all prompt evaluations in a stage (i.e., the paired setting) generally worsens performance. Examining the test performance, we observe that these conclusions generalize (Figure 7(b)).
@@
Figure 7: Normalized error (log scale) of the best prompt found by each HB validation instances sampling variant, averaged over benchmarks. Lower is better. Ribbons represent SE.
To test whether the design decisions of using supersets and using the same random validation instances for each prompt evaluation within a stage make a difference for the final performance of HB, we again conduct a linear mixed effects model analysis. We model the unaggregated performance at a fraction of $1.00$ total LLM calls involving random intercepts for each benchmark scenario (benchmark task and LLM combination). To test the global hypothesis that there is an effect of the design choices on performance, we test an intercept model against a model including an effect of the factor method (corresponding to an ablation variant). We test at the conservative $\alpha=0.01$ level.
For the validation error at a fraction of $1.00$, we reject the global null hypothesis of no effect of methods ($\chi^{2}(3)=730.85,p<1\mathrm{e}{-4}$). The pairwise results are:
@@
Examining test error at a fraction of $1.00$ we also reject the global null hypothesis of no effect of methods ($\chi^{2}(3)=207.12,p<1\mathrm{e}{-4}$). The pairwise results are:
@@
  • HB (no supersets, random instances for each prompt) vs. HB, $z=11.02,p<1\mathrm{e}{-4}$
@@
Summarizing, our results confirm that the design decisions made to adapt HB to prompt selection are effective: (1) The incumbent should be selected as the prompt with the lowest validation error among all prompts evaluated on the highest fidelity level. (2) Validation instances used to evaluate prompts in higher stages of a given bracket should be supersets of the validation instances used in lower stages. (3) Using the same (random) validation instances to evaluate prompts in each stage in general is beneficial compared to using truly random validation instances for each prompt, albeit this effect is comparably small.